# Pick up the Pace: A Parameter-Free Optimizer for Lifelong Reinforcement Learning

### v1

@@ 5 Discussion
Related work Combating loss of plasticity has been studied extensively in lifelong RL. A typical challenge for existing solutions is the tuning of their hyperparameters, which requires prior knowledge on the nature of the distribution shift, e.g., (Asadi et al., 2023; Ben-Iwhiwhu et al., 2023; Nikishin et al., 2023; Sokar et al., 2023; Mesbahi et al., 2024). An architectural modification called CReLU is studied in (Abbas et al., 2023), but our experiments suggest that its benefit might be specific to the Atari setup. Besides, Abel et al. (2018a, b) presented a theoretical analysis of skill transfer in lifelong RL, based on value iteration. Moreover, related contributions in nonstationary RL, where reward and state transition functions also change unpredictably, are limited to theoretical sequential decision-making settings with a focus on establishing complexity bounds (Roy et al., 2019; Cheung et al., 2020; Wei & Luo, 2021; Mao et al., 2022).
Our algorithm Pace builds on a long line of works on parameter-free OCO (see Section 3). To our knowledge, the only existing work applying parameter-free OCO to RL is (Jacobsen & Chan, 2021), which focuses on estimating of the value function (i.e. policy evaluation). Our scope is different, focusing on empirical RL in lifelong problems by exploring the key connection between parameter-free OCO and regularization.
Particularly, we are inspired by the Mechanic algorithm from (Cutkosky et al., 2023), which goes beyond the traditional convex setting of parameter-free OCO to handle stationary deep learning optimization tasks. Lifelong reinforcement learning, however, introduces a layer of complexity with its inherent nonstationarity. Furthermore, compared to Mechanic, Pace improves the scale tuner there (which is based on the coin-betting framework; Orabona & Pál, 2016) by the $\mathrm{erfi}$ algorithm that enjoys a better OCO performance guarantee. As an ablation study, we empirically compare Pace and Mechanic in the appendix (Figure 15). We find that Pace is slightly better, but both algorithms can mitigate the loss of plasticity, suggesting the effectiveness of the general “parameter-free” principle in lifelong RL.
@@
In general, the reference point $\theta_{\text{ref}}$ should be good or “safe” for Pace to perform effectively. One might presume that achieving this requires “warmstarting”, or pre-training using the underlying Base optimizer. While our experiments validate that such warmstarting is indeed beneficial (Appendix A), our main experiments show that even a random initialization of the policy’s weight serves as a good enough $\theta_{\text{ref}}$, even when tasks are similar (Figure 4).
@@
The success of Pace suggests that an adaptive form of regularization—anchoring to the reference point $\theta_{\mathrm{ref}}$—may suffice to counteract both mild and extreme forms of loss of plasticity. From this angle, we further elaborate the limitation of the $L_{2}$ regularization approach considered in (Kumar et al., 2023). It requires selecting a regularization strength parameter $\lambda$ through cross-validation, which is incompatible with the one-shot nature of lifelong learning settings. Furthermore, it is nontrivial to select the search grid: for example, we tried the $\lambda$-grid suggested by (Kumar et al., 2023), and there is no effective $\lambda$ value within the grid for the lifelong RL environments we consider. All the values are too small.
@@
In contrast, Pace offers a parameter-free solution that adapts dynamically with the data in an online manner. The scaling output of Pace adjusts autonomously to the ongoing conditions, consistently competing with well-tuned $\lambda$ values in the various environments, as demonstrated in the reward plots for CartPole, Acrobot, and LunarLander (Figure 7).
Figure 7: For each Gym Control environment and the initial ten tasks, we identified the best $\lambda$, which is the regularization strength that maximizes reward for each task’s specific distribution shift. Additionally we determined the best overall (well-tuned) $\lambda$ for each environment, optimizing performance across all tasks (CartPole $\lambda=25$, Acrobot $\lambda=15$, LunarLander $\lambda=45$). The results demonstrate that each environment and each task’s distribution shift is sensitive to different $\lambda$ and that Pace PPO performs competitively with each environment’s well-tuned regularization strength.
@@
Our results demonstrate the rapid adaptation of Pace, in lifelong RL problems with complicated function approximation. From the perspective of optimization, the latter requires tackling both nonconvexity and nonstationarity, which is typically regarded intractable in theory. Perhaps surprisingly, when approaching this complex problem using the theoretical insights from OCO, we observe compelling results. This suggests a certain “hidden convexity” in this problem, which could be an exciting direction for both theoretical and empirical research (e.g., policy gradient methods provably converge to global optimizers in linear quadratic control (Hu et al., 2023)).
@@
While Pace offers robust adaptability in nonstationary environments, it can exhibit suboptimal performance at the outset due to the time it requires to “warm up”. In the early stages of deployment, Pace might underperform compared to the baseline optimizer. We address this by proposing a warmstarting solution detailed in Appendix A, which helps bridge the initial performance gap. Currently, our results demonstrate the effectiveness of Pace in conjunction with PPO. However, Pace is fundamentally an optimizer that could broadly apply to other RL algorithms. Future work should extend the application of Pace to additional algorithms such as Soft Actor-Critic (Haarnoja et al., 2018) and Deep Q-Networks (Mnih et al., 2013).
@@ 6 Conclusion
In this work, we introduced Pace, a parameter-free optimizer for lifelong RL that leverages the principles of OCO. Our approach dynamically refines regularization in a data-dependent manner, eliminating the need for hyperparameter tuning. Through extensive experimentation in Procgen, Atari, and Gym Control environments, we demonstrated that Pace effectively mitigates loss of plasticity and rapidly adapts to new distribution shifts, where baseline methods fail. Pace’s success leads to a compelling takeaway: empirical lifelong RL scenarios may exhibit more convex properties than previously appreciated, and might inherently benefit from parameter-free OCO approaches.
Therefore, in lifelong RL, to rapidly adapt to new changes without prior knowledge, we suggest you “Pick up (the) Pace”.
@@ 7 Acknowledgments
Appendix A Warmstarting
@@
In our theoretical framework, we hypothesize that a robust parameter initialization, denoted as $\theta_{\mathrm{ref}}$, could enhance the performance of our models, suggesting that empirical implementations might benefit from initializing parameters using a base optimizer such as Adam prior to deploying Pace. Contrary to this assumption, our experimental results detailed in Section 4 reveal that warmstarting is not essential for Pace’s success. Below, we examine the performance of Adam PPO and Pace PPO when warmstarted.
@@
Both Pace PPO and Adam PPO were warmstarted using Adam for the initial 150,000 steps in all games for the Atari and Procgen environments, and for the first 30 steps in the Gym Control experiments. As seen in Figure 8, in games like Starpilot, Fruitbot, and Dodgeball, Pace PPO surpasses Adam PPO in the first level/task of the online setup, with its performance closely matching that of Adam PPO in Chaser. Importantly, Pace PPO continues to circumvent the loss of plasticity encountered by Adam PPO, even when both are warmstarted. This makes sense since all of the distributions share some foundational game dynamics; the initial learning phases likely explore these dynamics, so leveraging a good parameter initialization to regularize in this early region can be beneficial for Pace—we observe that forward transfer occurs somewhat in later level distribution shifts as the reward does not drop back to zero where it initially started from.
@@
Our findings indicate that warmstarting does not confer a significant advantage in the Atari games. This makes sense because a parameter initialization that is good in one game setting is likely a random parameterization for another setting, which is equivalent to the setup without warmstarting where Pace regularizes towards a random parameter initialization. In the Gym Control experiments although warmstarted Pace PPO manages to avoid the extreme plasticity loss and policy collapse seen in warmstarted Adam PPO, it does not perform as well as non-warmstarted Pace PPO. This variability underscores that the efficacy of warmstarting is environment-specific and highlights the challenge in predicting when Adam PPO may achieve a parameter initialization that is advantageous for Pace PPO to regularize towards.
@@
From an overall perspective, warmstarting Pace PPO in every setting still shows substantial improvement over Adam PPO (Table 2)
@@
Figure 8: Reward in the lifelong Procgen environments for StarPilot, Dodgeball, Fruitbot, and Chaser with warmstarted Pace PPO and warmstarted Adam PPO. Inital performance of Pace PPO is improved with warmstarting and continues to avoid loss of plasticity.
@@
Figure 9: Reward in the lifelong Atari environments with warmstarted Pace PPO and warmstarted Adam PPO. No significant benefit is found by warmstarting Pace compared to not warmstarting it.
@@
Figure 10: Reward in the lifelong Gym Control environments for CartPole-v1, Acrobot-v1, and LunarLander-v2 with warmstarted Pace PPO and warmstarted Adam PPO.
@@
Table 2: Cumulative sum of mean episode reward over all distributions for Adam PPO warmstarted and Pace PPO warmstarted on Procgen, Atari, and Gym Control environments. Rewards are scaled by $10^{5}$; higher is better.
Environment Adam PPO Pace PPO (Ours)
@@
Appendix B Gravity Based Distribution Shifts
@@
Appendix C Scaling-Value Convergence
As discussed in the algorithm section (see Section 3), Pace operates as a meta-algorithm on top of a standard optimizer, denoted as Base. The crucial component of Pace involves the dynamic adjustment of the scaling parameter $S_{t+1}$, managed by the tuner algorithm (Algorithm 2). This parameter is data-dependent and typically ranges between $[0,1]$. The weight update $\theta_{t+1}$ is consequently defined as a convex combination of the current optimizer’s weight $\theta_{t}^{\textsc{Base}}$ and a predetermined reference point $\theta_{\text{ref}}$.
@@ Appendix C Scaling-Value Convergence
The convergence of the scaling parameter $S_{t+1}$ observed across the Procgen and Gym Control environments, as depicted in Figures 12 and 14, reflects a good scaling value that effectively determines the strength of regularization towards the initialization points, yielding robust empirical outcomes in lifelong RL settings. Interestingly, in Procgen environments, this converged scaling value exhibits consistency across various games, typically hovering between 0.02 and 0.03, as shown in Figure 12. In contrast, in the Gym Control environments, the scaling values are lower, ranging between 0.005 and 0.01, as illustrated in Figure 14.
Appendix D Comparison to Mechanic
In our analysis, we extend the examination to other OCO-based optimizers within the lifelong RL framework. Figure 15 presents a comparative assessment of Pace PPO and Mechanic PPO (Cutkosky et al., 2023) in the Starpilot game from the Procgen suite. The results indicate that while Mechanic effectively mitigates plasticity loss and adapts quickly to new distribution shifts, it slightly underperforms in comparison to Pace. This suggests potential for further explorations and a broader evaluation of OCO-based optimizers in lifelong learning environments.
Figure 15: Reward in the lifelong StarPilot environment with both Pace PPO and Mechanic PPO. Mechanic PPO performs similarly to Pace PPO, although slightly underperforming Pace PPO.
Appendix E Experimental Setup
@@
Pace
Pace, for all experiments, was implemented using the same experiment-specific baseline architectures and baseline optimizer. For the Procgen and Atari experiments, the base Adam optimizer was configured as the same as baseline, with a learning rate of 0.001, and for the Gym Control experiments, a learning rate of 0.01 was used. Both learning rates were tested for all experiments and found to have negligible differences in performance outcomes. Other than the learning rate, we use the default Adam parameters, including weight decay and betas, followed by the specifications outlined in the PyTorch Documentation. [^4] https://pytorch.org/docs/stable/generated/torch.optim.Adam.html
The setup for Pace included $\beta$ values for adaptive gradient adjustments: 0.9, 0.99, 0.999, 0.9999, 0.99999, and 0.999999. Both $S_{t}$ and $\varepsilon$ were initially set to ($1\times 10^{-8}$). Modifications were made to a PyTorch error function library, which accepts complex inputs to accommodate the necessary computations for the imaginary error function. This library can be found at Torch Erf GitHub. [^5] https://github.com/redsnic/torch_erf
@@
The Procgen and Atari experiments were conducted with 8 seeds/runs, while the Gym Control experiments utilized 25 seeds/runs. The exception was in the $L_{2}$ initialization experiments, which used 15 seeds/runs per regularization strength. In Figures 4, 5, 6, 7, 8, 9, 10, 11, the plotted lines represent the mean of all of the mean episode rewards from the different seeds/runs, and the shaded error bands indicate the standard deviation of all of the mean episode rewards from the different seeds/runs.

### v2

@@ 5 Discussion
Related work Combating loss of plasticity has been studied extensively in lifelong RL. A typical challenge for existing solutions is the tuning of their hyperparameters, which requires prior knowledge on the nature of the distribution shift, e.g., (Asadi et al., 2024; Ben-Iwhiwhu et al., 2022; Nikishin et al., 2024; Sokar et al., 2023; Mesbahi et al., 2024). An architectural modification called CReLU is studied in (Abbas et al., 2023), but our experiments suggest that its benefit might be specific to the Atari setup. Besides, Abel et al. (2018a, b) presented a theoretical analysis of skill transfer in lifelong RL, based on value iteration. Moreover, related contributions in nonstationary RL, where reward and state transition functions also change unpredictably, are limited to theoretical sequential decision-making settings with a focus on establishing complexity bounds (Roy et al., 2019; Cheung et al., 2020; Wei & Luo, 2021; Mao et al., 2020).
Our algorithm Trac builds on a long line of works on parameter-free OCO (see Section 3). To our knowledge, the only existing work applying parameter-free OCO to RL is (Jacobsen & Chan, 2021), which focuses on estimating the value function (i.e., policy evaluation). Our scope is different, focusing on empirical RL in lifelong problems by exploring the key connection between parameter-free OCO and regularization.
Particularly, we are inspired by the Mechanic algorithm from (Cutkosky et al., 2023), which goes beyond the traditional convex setting of parameter-free OCO to handle stationary deep learning optimization tasks. Lifelong reinforcement learning, however, introduces a layer of complexity with its inherent nonstationarity. Furthermore, compared to Mechanic, Trac improves the scale tuner there (which is based on the coin-betting framework; Orabona & Pál, 2016) by the $\mathrm{erfi}$ algorithm that enjoys a better OCO performance guarantee. As an ablation study, we empirically compare Trac and Mechanic in the Appendix E (Figure 16). We find that Trac is slightly better, but both algorithms can mitigate the loss of plasticity, suggesting the effectiveness of the general “parameter-free” principle in lifelong RL.
Trac encourages positive transfer
In our experiments, we observe that Trac’s reward decline due to distribution shifts is less severe than that of baseline methods. These results may suggest that Trac facilitates positive transfer between related tasks in a lifelong learning setting. To investigate this further, we compared Trac to a privileged weight-resetting approach, where the network’s parameters are reset for each new task, in the Gym Control environments (see Appendix A). Our results show that Trac maintains higher rewards during tasks than privileged weight-resetting and avoids declining to the same low reward levels as privileged weight-resetting at the start of a new task (Figure 8).
@@
In general, the reference point $\theta_{\text{ref}}$ should be good or “safe” for Trac to perform effectively. One might presume that achieving this requires “warmstarting”, or pre-training using the underlying Base optimizer. While our experiments validate that such warmstarting is indeed beneficial (Appendix B), our main experiments show that even a random initialization of the policy’s weight serves as a good enough $\theta_{\text{ref}}$, even when tasks are similar (Figure 4).
@@
The success of Trac suggests that an adaptive form of regularization—anchoring to the reference point $\theta_{\mathrm{ref}}$—may suffice to counteract both mild and extreme forms of loss of plasticity. From this angle, we further elaborate the limitation of the $L_{2}$ regularization approach considered in (Kumar et al., 2023). It requires selecting a regularization strength parameter $\lambda$ through cross-validation, which is incompatible with the one-shot nature of lifelong learning settings. Furthermore, it is nontrivial to select the search grid: for example, we tried the $\lambda$-grid suggested by (Kumar et al., 2023), and there is no effective $\lambda$ value within the grid for the lifelong RL environments we consider. All the values are too small.
@@
In contrast, Trac offers a parameter-free solution that adapts dynamically with the data in an online manner. The scaling output of Trac adjusts autonomously to the ongoing conditions, consistently competing with well-tuned $\lambda$ values in the various environments, as demonstrated in the reward plots for CartPole, Acrobot, and LunarLander (Figure 7).
Figure 7: For each Gym Control environment and the initial ten tasks, we identified the best $\lambda$, which is the regularization strength that maximizes reward for each task’s specific distribution shift. Additionally we determined the best overall (well-tuned) $\lambda$ for each environment, optimizing performance across all tasks (CartPole $\lambda=25$, Acrobot $\lambda=15$, LunarLander $\lambda=45$). The results demonstrate that each environment and each task’s distribution shift is sensitive to different $\lambda$ and that Trac PPO performs competitively with each environment’s well-tuned regularization strength.
@@
Our results demonstrate the rapid adaptation of Trac, in lifelong RL problems with complicated function approximation. From the perspective of optimization, the latter requires tackling both nonconvexity and nonstationarity, which is typically regarded intractable in theory. Perhaps surprisingly, when approaching this complex problem using the theoretical insights from OCO, we observe compelling results. This suggests a certain “hidden convexity” in this problem, which could be an exciting direction for both theoretical and empirical research (e.g., policy gradient methods provably converge to global optimizers in linear quadratic control (Hu et al., 2023)).
@@
While Trac offers robust adaptability in nonstationary environments, it can exhibit suboptimal performance at the outset due to the time it requires to “warm up”. In the early stages of deployment, Trac might underperform compared to the baseline optimizer. We address this by proposing a warmstarting solution detailed in Appendix B, which helps bridge the initial performance gap. Currently, our results demonstrate the effectiveness of Trac in conjunction with PPO. However, Trac is fundamentally an optimizer that could broadly apply to other RL algorithms. Future work should extend the application of Trac to additional algorithms such as Soft Actor-Critic (Haarnoja et al., 2018) and Deep Q-Networks (Mnih et al., 2015).
@@ 6 Conclusion
In this work, we introduced Trac, a parameter-free optimizer for lifelong RL that leverages the principles of OCO. Our approach dynamically refines regularization in a data-dependent manner, eliminating the need for hyperparameter tuning. Through extensive experimentation in Procgen, Atari, and Gym Control environments, we demonstrated that Trac effectively mitigates loss of plasticity and rapidly adapts to new distribution shifts, where baseline methods fail. Trac’s success leads to a compelling takeaway: empirical lifelong RL scenarios may exhibit more convex properties than previously appreciated, and might inherently benefit from parameter-free OCO approaches.
@@ 7 Acknowledgments
Appendix A Trac Encourages Positive Transfer
To explore whether Trac encourages positive transfer, we introduce a privileged weight-reset baseline. This baseline is "privileged" in the sense that it knows when a distribution shift is introduced and resets the parameters to a random initialization at the start of each new task. We applied this baseline to three Gym control tasks: CartPole-v1, Acrobot-v1, and LunarLander-v2, and compared it to Trac PPO and Adam PPO, as shown in Figure 8.
We observe that the privileged weight-reset baseline exhibits spikes in reward at the beginning of each new task. Surprisingly, Trac maintains even higher rewards than the privileged weight-reset baseline, even at its peak learning phases. Additionally, Trac’s reward does not decline to the reward seen at the start of new tasks with privileged weight-resetting (Trac does not have to "start over" with each task), suggesting that Trac successfully transfers skills positively between tasks.
Figure 8: Reward comparison of Trac PPO, Adam PPO, and privileged weight-resetting on Cartpole-v1, Acrobot-v1, and LunarLander-v2. Trac PPO encourages positive transfer between tasks.
Appendix B Warmstarting
@@
In our theoretical framework, we hypothesize that a robust parameter initialization, denoted as $\theta_{\mathrm{ref}}$, could enhance the performance of our models, suggesting that empirical implementations might benefit from initializing parameters using a base optimizer such as Adam prior to deploying Trac. Contrary to this assumption, our experimental results detailed in Section 4 reveal that warmstarting is not essential for Trac’s success. Below, we examine the performance of Adam PPO and Trac PPO when warmstarted.
@@
Both Trac PPO and Adam PPO were warmstarted using Adam for the initial 150,000 steps in all games for the Atari and Procgen environments, and for the first 30 steps in the Gym Control experiments. As seen in Figure 9, in games like Starpilot, Fruitbot, and Dodgeball, Trac PPO surpasses Adam PPO in the first level/task of the online setup, with its performance closely matching that of Adam PPO in Chaser. Importantly, Trac PPO continues to avoid the loss of plasticity encountered by Adam PPO, even when both are warmstarted. This makes sense since all of the distributions share some foundational game dynamics; the initial learning phases likely explore these dynamics, so leveraging a good parameter initialization to regularize in this early region can be beneficial for Trac—we observe that forward transfer occurs somewhat in later level distribution shifts as the reward does not drop back to zero where it initially started from.
@@
Our findings indicate that warmstarting does not confer a significant advantage in the Atari games. This makes sense because a parameter initialization that is good in one game setting is likely a random parameterization for another setting, which is equivalent to the setup without warmstarting where Trac regularizes towards a random parameter initialization. In the Gym Control experiments although warmstarted Trac PPO manages to avoid the extreme plasticity loss and policy collapse seen in warmstarted Adam PPO, it does not perform as well as non-warmstarted Trac PPO. This result underscores that the efficacy of warmstarting is environment-specific and highlights the challenge in predicting when Adam PPO may achieve a parameter initialization that is advantageous for Trac PPO to regularize towards.
@@
From an overall perspective, warmstarting Trac PPO in every setting still shows substantial improvement over Adam PPO (Table 2).
@@
Figure 9: Reward in the lifelong Procgen environments for StarPilot, Dodgeball, Fruitbot, and Chaser with warmstarted Trac PPO and warmstarted Adam PPO. Inital performance of Trac PPO is improved with warmstarting and continues to avoid loss of plasticity.
@@
Figure 10: Reward in the lifelong Atari environments with warmstarted Trac PPO and warmstarted Adam PPO. No significant benefit is found by warmstarting Trac PPO compared to not warmstarting it.
@@
Figure 11: Reward in the lifelong Gym Control environments for CartPole-v1, Acrobot-v1, and LunarLander-v2 with warmstarted Trac PPO and warmstarted Adam PPO. Trac PPO still avoids loss of plasticity and policy collapse.
@@
Table 2: Cumulative sum of mean episode reward over all distributions for Adam PPO warmstarted and Trac PPO warmstarted on Procgen, Atari, and Gym Control environments. Rewards are scaled by $10^{5}$; higher is better.
Environment Adam PPO Trac PPO (Ours)
@@
Appendix C Gravity Based Distribution Shifts
@@
Appendix D Scaling-Value Convergence
As discussed in the algorithm section (see Section 3), Trac operates as a meta-algorithm on top of a standard optimizer, denoted as Base. The crucial component of Trac involves the dynamic adjustment of the scaling parameter $S_{t+1}$, managed by the tuner algorithm (Algorithm 2). This parameter is data-dependent and typically ranges between $[0,1]$. The weight update $\theta_{t+1}$ is consequently defined as a convex combination of the current optimizer’s weight $\theta_{t}^{\textsc{Base}}$ and a predetermined reference point $\theta_{\text{ref}}$.
@@ Appendix D Scaling-Value Convergence
The convergence of the scaling parameter $S_{t+1}$ observed across the Procgen and Gym Control environments, as depicted in Figures 13 and 15, reflects a good scaling value that effectively determines the strength of regularization towards the initialization points, yielding robust empirical outcomes in lifelong RL settings. Interestingly, in Procgen environments, this converged scaling value exhibits consistency across various games, typically hovering between 0.02 and 0.03, as shown in Figure 13. In contrast, in the Gym Control environments, the scaling values are lower, ranging between 0.005 and 0.01, as illustrated in Figure 15.
Appendix E Comparison to Mechanic
In our analysis, we extend the examination to other OCO-based optimizers within the lifelong RL framework. Figure 16 presents a comparative assessment of Trac PPO and Mechanic PPO (Cutkosky et al., 2023) for Procgen and Gym Control tasks. The results indicate that while Mechanic effectively mitigates plasticity loss and adapts quickly to new distribution shifts, it slightly underperforms in comparison to Trac. This suggests potential for further explorations and a broader evaluation of OCO-based optimizers in lifelong learning environments.
Figure 16: Reward in the Procgen and Control environments with both Trac PPO and Mechanic PPO. Mechanic PPO performs similarly to Trac PPO, although slightly underperforming Trac PPO.
Appendix F Experimental Setup
@@
Trac
Trac, for all experiments, was implemented using the same experiment-specific baseline architectures and baseline optimizer. For the Procgen and Atari experiments, the base Adam optimizer was configured as the same as baseline, with a learning rate of 0.001, and for the Gym Control experiments, a learning rate of 0.01 was used. Both learning rates were tested for all experiments and found to have negligible differences in performance outcomes. Other than the learning rate, we use the default Adam parameters, including weight decay and betas, followed by the specifications outlined in the PyTorch Documentation. [^5] https://pytorch.org/docs/stable/generated/torch.optim.Adam.html
The setup for Trac included $\beta$ values for adaptive gradient adjustments: 0.9, 0.99, 0.999, 0.9999, 0.99999, and 0.999999. Both $S_{t}$ and $\varepsilon$ were initially set to ($1\times 10^{-8}$). Modifications were made to a PyTorch error function library, which accepts complex inputs to accommodate the necessary computations for the imaginary error function. This library can be found at Torch Erf GitHub. [^6] https://github.com/redsnic/torch_erf
@@
The Procgen and Atari experiments were conducted with 8 seeds/runs, while the Gym Control experiments utilized 25 seeds/runs. The exception was in the $L_{2}$ initialization experiments, which used 15 seeds/runs per regularization strength. In Figures 4, 5, 6, 7, 9, 10, 11, 12, the plotted lines represent the mean of all of the mean episode rewards from the different seeds/runs, and the shaded error bands indicate the standard deviation of all of the mean episode rewards from the different seeds/runs.